# It Is Giving Major Satisfaction: Why Fairness Matters for Software Practitioners

### v2

@@ 5.5. Moderation Analysis
In our study, we specifically examined the potential moderating effects of various demographic factors, as described in Section 6.1. Due to the imbalance in the distribution of demographic categories among our respondents (as shown in Table 4), initially, we were only able to effectively analyze the moderating effects for four demographic groups: gender, ethnicity representation, and work limitation. Subsequently, we simplified the categories for the remaining demographic groups to improve the analysis. For instance, ethnicity was grouped into Asian and Non-Asian categories, and work experience was divided into less than 3 years and more than 3 years. This refinement allowed us to effectively analyze the moderating effects for one additional demographic group: work experience.
However, the limited number of respondents in other remaining demographic categories, namely age, ethnicity, role and role nature restricted our ability to conduct reliable moderation analyses for these groups. This limitation is due to the potential for biased estimates caused by imbalanced distribution within the sample (Zhang and Wang, 2017).
@@ 6.2. Survey Responses Summary
Our survey examined software practitioners’ perceptions of fairness, job satisfaction, and their experiences of fairness in various SE contexts. The survey results indicate a mixed perception of fairness across different dimensions among software practitioners. While interpersonal fairness is generally perceived positively by 90% respondents, there seems to be moderate fairness with informational (79%), procedural (77%), and distributive fairness (77%).
@@ 6.3. RQ1: Does fairness perception shape job satisfaction among software practitioners? If so, how?
The results of the analysis are summarized in the first section of Table 6. We considered relationships to be impactful if they were statistically significant ($p<0.05$) and had OR exceeding 50%, to ensure that the identified factors not only demonstrate a reliable statistical relationship but also have a practically meaningful impact on job satisfaction (see (fig, 2024)). Our analysis demonstrates that perceptions of fairness across various dimensions significantly shape job satisfaction among software practitioners.
@@
For example, among interpersonal fairness items, being treated with respect (Int2) had the greatest effect, increasing the likelihood of higher overall job satisfaction by 240%. This suggests that when employees feel respected, they are more than twice as likely to report higher levels of job satisfaction. Similarly, within procedural fairness, ethical and moral processes (P1) increased the odds of higher overall job satisfaction by 128%, indicating that ethical practices are crucial for building trust between employees and management.
Additionally, accurate information (P3), another procedural fairness item, was shown to increase satisfaction with job security by 113%. This finding suggests that when software practitioners receive clear, reliable, and accurate information about organizational processes and decisions, they feel more secure in their roles.
@@
Table 6. Ordinal Regression Analysis Result
@@
D1 + + +
D2 + + + +
D3 +
D4 + + + +
P1 + + +
P2 + +
P3 + + + +
P4 + +
P5 +
P6 + +
P7 +
Int1 + + +
Int2 + + +
Int3 + + +
Int4 +
Inf1
Inf2 + +
Inf3 +
Inf4 +
Inf5 +
@@
C1
C2
C3
C4 + +
C5 + +
C6 +
C7 + + +
C8 +
C9 + + + + +
C10 + + +
@@
Inf1
Inf2
Inf3
@@ 6.5. RQ3: How do fairness perceptions influence job satisfaction in specific SE contexts?
Overall job satisfaction is significantly correlated with a high level of fairness experience in several specific SE contexts, particularly in the areas of demand (C10), authorship (C9), allocation of work (C7), and policy (C4). Additionally, as observed in RQ1, All four dimensions of fairness, distributive, procedural, interpersonal, and informational fairness are linked to higher overall satisfaction. This suggests that activities associated with all four fairness dimensions in these contexts contribute to higher overall job satisfaction.
Further analysis indicates that, as noted in RQ1, these four dimensions also play a crucial role in enhancing job security (SEC). In the context of RQ3, we found a positive and significant correlation in the contexts of authorship and evaluation. This means that software practitioners feel more secure in their jobs when, during evaluations such as performance reviews, or when their contributions, like coding solutions or innovative ideas, are properly acknowledged in project meetings or official documentation, all four fairness dimensions are fulfilled.
@@
Gender. As noted in RQ2, distributive, procedural and informational fairness were found to have the significant impact on female developers’ satisfaction with their manager (MNG). While in RQ3, we found that the contexts of policy (C4) and evaluation (C5) also significantly affects female developers’ satisfaction with their manager (MNG). This suggests that to increase female developers’ satisfaction with their managers, organizations should focus on ensuring fair outcomes, accurate information, unbiased processes, and tailored communication during policy implementation and evaluations.
@@
Ethnicity Representation. We found that unbiased processes (P4) and the ability to participate in decision-making (P5) are particularly important for underrepresented practitioners’ satisfaction with performance (PERF), especially in the contexts of policy (C4) and demand (C10). This suggests that underrepresented developers value having unbiased processes and the opportunity to participate in decision-making, especially when it comes to policies and work expectations. When these developers feel they can voice their opinions and see that decisions are made fairly, their satisfaction with their performance increases.
@@
Work Experience. In RQ2, we found that less experienced practitioners feel more satisfied with their performance when treated with politeness, dignity, and respect, and when inappropriate behavior or comments are absent. In the current analysis, satisfaction with performance (PRF) seems to be particularly influenced by their working hours (C6).
@@
Work Limitation. The analysis in RQ2 also shows that software practitioners with work limitations feel significantly stronger about their satisfaction with compensation (COM), particularly when it comes to being treated politely (Int1), with dignity (Int3), avoiding inappropriate comments or behavior (Int4), and receiving comprehensive explanations (Inf5). In the current analysis, we see that the context of demand (C10) affects practitioners with work limitations more strongly than those without when it comes to their satisfaction with compensation (COM).
@@
Takeaway 4: Regarding job satisfaction, female practitioners are more impacted by fairness in policy implementation and evaluations. Underrepresented practitioners value fairness in policy implementation and during high-demand situations. Less experienced practitioners place greater emphasis on fairness when navigating working hours, while practitioners with work limitations prioritize fairness during demanding situations.
@@ 8. Threats to Validity
Another potential threat is the imbalance in the distribution of certain demographic categories, which limited our ability to conduct reliable moderation analyses across all groups. To mitigate this, we focused our moderation analyses on demographic groups with sufficient data: gender, ethnicity representation, work limitation, and work experience. By concentrating on these groups, we ensured that the moderation analyses were based on more robust sample sizes, reducing the likelihood of bias in our findings. Although we were unable to analyze other categories, such as age, ethnicity, role, and role nature, our approach ensures that the results we report are based on well-supported data and that we minimized the risk of skewed conclusions.

### v3

@@ 4.5.2. Moderation Analysis
In our study, we specifically examined the potential moderating effects of various demographic factors, as described in Section 5.1. Due to the imbalance in the distribution of demographic categories among our respondents (as shown in Table 4), initially, we were only able to effectively analyze the moderating effects for three demographic groups: gender, ethnicity representation, and work limitation. Subsequently, we simplified the categories for the remaining demographic groups to improve the analysis. For instance, ethnicity was grouped into Asian and Non-Asian categories, and work experience was divided into less than 3 years and more than 3 years. This refinement allowed us to effectively analyze the moderating effects for one additional demographic group: work experience.
However, the limited number of respondents in other remaining demographic categories, namely age, ethnicity, role and role nature restricted our ability to conduct reliable moderation analyses for these groups. This limitation is due to the potential for biased estimates caused by imbalanced distribution within the sample (Zhang and Wang, 2017; Griffin et al., 2023).
For RQ2, we sought to identify whether these demographic characteristics alter the impact of fairness perceptions on job satisfaction, and if so, in what ways. For instance, we hypothesized that certain demographic groups, such as those with less representation in the workplace, might experience a stronger relationship between biased processes (part of procedural fairness) and their satisfaction with job security (SEC) compared to more represented groups.
For RQ3.1, we aimed to examine how fairness perception influencing job satisfaction differ across demographic groups within the SE context. For instance, if we observe that being treated politely, with respect, and dignity during evaluations significantly impacts job satisfaction, it would be important to explore which demographic groups are most affected by this factor. This could suggest that certain groups, such as female developers, find this aspect more impactful, indicating that gender may play a key role in shaping this relationship.
@@ 5.2. Survey Responses Summary
Our survey examined software practitioners’ perceptions of fairness, job satisfaction, and their experiences of fairness in various SE contexts. The survey results (Table LABEL:tab:itemlevel) indicate a mixed perception of fairness across different dimensions among software practitioners. While interpersonal fairness is generally perceived positively by 90% respondents, there seems to be moderate fairness with informational (79%), procedural (77%), and distributive fairness (77%).
It is worth noting that while interpersonal fairness received generally positive responses (median and mean = 4), the other dimensions tended to cluster around a median and mean of 3, indicating a more neutral stance. This neutrality could mean several points. It might suggest that respondents view the situation as fair enough and see no major problems, especially if they have not encountered obvious unfairness. But it could also point to uncertainty or disengagement. Some respondents may not have thought much about fairness in their workplace or may lack opportunities to reflect on or discuss it openly.
@@ 5.3. RQ1: Does fairness perception shape job satisfaction among software practitioners? If so, how?
The results of the analysis are summarized in the first section of Table 6. We considered relationships to be impactful if they were statistically significant ($p<0.05$) and had OR exceeding 50%, to ensure that the identified factors not only demonstrate a reliable statistical relationship but also have a practically meaningful impact on job satisfaction (see (fig, 2024)). We express OR as percentage increases in the likelihood of higher job satisfaction associated with stronger perceptions of fairness. Our analysis demonstrates that perceptions of fairness across various dimensions significantly shape job satisfaction among software practitioners.
@@
For example, among interpersonal fairness items, being treated with respect (Int2) had the greatest effect, increasing the likelihood of higher overall job satisfaction by 240% (OR = 3.40, 95% CI[2.19, 5.37]). This suggests that when employees feel respected, they are more than twice as likely to report higher levels of job satisfaction. Similarly, within procedural fairness, ethical and moral processes (P1) increased the odds of higher overall job satisfaction by 128% (OR = 2.28, 95% CI[1.54, 3.37]), indicating that ethical practices are crucial for building trust between employees and management.
Additionally, accurate information (P3), another procedural fairness item, was shown to increase satisfaction with job security by 113% (OR = 2.13, 95% CI[1.43, 3.16]). This finding suggests that when software practitioners receive clear, reliable, and accurate information about organizational processes and decisions, they feel more secure in their roles.
@@
Table 6. Ordinal Regression Analysis Result. Abbreviations in the first row refer to job satisfaction composite factors and are explained in Table 3.
@@
D1: Outcome is justified by the performance + + +
D2: Outcome is appropriate for the completed work + + + +
D3: Outcome reflecting contribution +
D4: Outcome reflecting effort + + + +
P1: Ethical and moral processes + + +
P2: Voicing opinions + +
P3: Accurate information + + + +
P4: Free of bias processes + +
P5: Have a say in making decisions +
P6: Consistency in applying processes + +
P7: Appealing decisions +
Int1: Treated politely + + +
Int2: Treated with respect + + +
Int3: Treated with dignity + + +
Int4: No inappropriate comments or behaviour +
Inf1: Direct communication
Inf2: Reasonable explanations + +
Inf3: Timely communication +
Inf4: Tailored communication +
Inf5: Comprehensive explanations +
@@
C1: Customer/client Relations
C2: Recruitment
C3: Treatment
C4: Policy + +
C5: Evaluation + +
C6: Working Hours +
C7: Allocation of Work + + +
C8: Income +
C9: Authorship + + + + +
C10: Demand + + +
@@
Inf1: Direct communication
Inf2: Reasonable explanations
Inf3: Timely communication
@@ 5.5. RQ3: How do fairness perceptions influence job satisfaction in specific SE contexts?
Overall job satisfaction is significantly correlated with a high level of fairness experience in several specific SE contexts, particularly in the areas of demand (C10) (OR = 1.73, 95% CI[1.27, 2.37]), authorship (C9) (OR = 1.59, 95% CI[1.15, 2.18]), allocation of work (C7) (OR = 1.56, 95% CI[1.13, 2.17]), and policy (C4) (OR = 1.54, 95% CI[1.14, 2.08]). Additionally, as observed in RQ1, All four dimensions of fairness, distributive, procedural, interpersonal, and informational fairness are linked to higher overall satisfaction. This suggests that activities associated with all four fairness dimensions in these contexts contribute to higher overall job satisfaction.
Further analysis indicates that, as noted in RQ1, these four dimensions also play a crucial role in enhancing job security (SEC). In the context of RQ3, we found a positive and significant correlation in the contexts of authorship (OR = 1.77, 95% CI[1.29, 2.42]) and evaluation (OR = 1.77, 95% CI[1.30, 2.41]). This means that software practitioners feel more secure in their jobs when, during evaluations such as performance reviews, or when their contributions, like coding solutions or innovative ideas, are properly acknowledged in project meetings or official documentation, all four fairness dimensions are fulfilled.
@@
Gender. As noted in RQ2, distributive, procedural and informational fairness were found to have the significant impact on female developers’ satisfaction with their manager (MNG). While in RQ3.1, we found that the contexts of policy (C4) (OR = 0.34, 95% CI[0.13, 0.77]) and evaluation (C5) (OR = 0.40, 95% CI[0.17, 0.89]) also significantly affects female developers’ satisfaction with their manager (MNG). This suggests that to increase female developers’ satisfaction with their managers, organizations should focus on ensuring fair outcomes, accurate information, unbiased processes, and tailored communication during policy implementation and evaluations.
@@
Ethnicity Representation. We found that unbiased processes (P4) and consistent processes (P6) are particularly important for underrepresented practitioners’ satisfaction with performance (PERF), especially in the contexts of policy (C4) (OR = 3.61, 95% CI[1.34, 8.91]) and allocation of work (C7) (OR = 3.08, 95% CI[1.14, 7.77]). Our findings suggest that for underrepresented developers, having unbiased and consistent processes is especially important for their satisfaction with performance, particularly in the context of policy decisions and work allocation. When these processes are in place, it appears to support a stronger sense of fairness, which may contribute to greater performance satisfaction.
@@
Work Experience. In RQ2, we found that less experienced practitioners feel more satisfied with their performance when treated with politeness, dignity, and respect, and when inappropriate behavior or comments are absent. In the current analysis, satisfaction with performance (PRF) seems to be particularly influenced by their working hours (C6) (OR = 0.43, 95% CI[0.21, 0.89]).
@@
Work Limitation. The analysis in RQ2 also shows that software practitioners with work limitations feel significantly stronger about their satisfaction with compensation (COM), particularly when it comes to being treated politely (Int1), with dignity (Int3), avoiding inappropriate comments or behavior (Int4), and receiving comprehensive explanations (Inf5). In the current analysis, we see that the context of demand (C10) affects practitioners with work limitations more strongly than those without when it comes to their satisfaction with compensation (COM) (OR = 2.92, 95% CI[1.24, 6.83]).
@@
Takeaway 4: Regarding job satisfaction, female practitioners are more impacted by fairness in policy implementation and evaluations. Underrepresented practitioners value fairness in policy implementation and during work allocation. Less experienced practitioners place greater emphasis on fairness when navigating working hours, while practitioners with work limitations prioritize fairness during demanding situations.
@@ 7. Threats to Validity
Like many survey-based studies, our findings rely on self-reported responses, which may be subject to biases such as social desirability or recall bias. While we mitigated some of these risks by ensuring anonymity and using validated survey instruments, we recognize that participants may still over- or under-report certain experiences. Future studies could strengthen construct validity by combining self-reported data with qualitative methods such as interviews or focus groups to triangulate findings and provide deeper insights.
Another construct-related concern is whether all respondents were legitimate software practitioners. Since the survey was anonymous and not linked to participants’ online profiles, we could not verify professional identity. However, given the absence of monetary incentives, and the use of recruitment channels (e.g., LinkedIn, SE workshops), we consider the risk of identity falsification to be minimal. Respondents had no clear motivation to misrepresent their role, and the targeted nature of our outreach makes it unlikely that the survey reached individuals without relevant SE experience.
@@
Another potential threat is the imbalance in the distribution of certain demographic categories, which limited our ability to conduct reliable moderation analyses across all groups. To mitigate this, we focused our analyses on demographic variables with sufficiently large subgroup sizes to ensure statistical reliability and reduce the risk of biased estimates, as recommended in prior research (Zhang and Wang, 2017; Griffin et al., 2023). Specifically, we analyzed gender, ethnicity representation, work limitation, and work experience, where each category included an adequate number of respondents. Although we were unable to analyze other categories, such as age, ethnicity, role, and role nature, this approach ensured that our moderation analyses were grounded in robust data, and to minimize the likelihood of skewed conclusions due to small sample sizes. Further clarification on this methodological decision is available in Section 4.5.2.
@@
Finally, while our sample is not representative in a statistical sense, we ensured heterogeneity by recruiting participants across gender, age, team composition, roles, experience levels. This diversity improves the credibility of our findings in line with SE research norms for theory-building studies (Baltes and Ralph, 2022). We acknowledge that our findings may not generalize to all software practitioners globally, especially those outside our sample’s distribution. As Baltes and Ralph caution (Baltes and Ralph, 2022), non-probability samples may limit statistical generalization. We address this by clearly defining our target population and transparently reporting recruitment procedures. Future studies with larger and more representative samples can build on our findings to deepen and extend this line of research.